OpenStudy (anonymous):

Which value of z makes the equation true? z – 4.7 = –0.9 z =

OpenStudy (anonymous):

@BPDlkeme234 pls help

OpenStudy (anonymous):

@nechirwan

OpenStudy (anonymous):

add 4.7 to the other side to -.9 to find the value of x

OpenStudy (anonymous):

x = 5.6

OpenStudy (anonymous):

pls help @BPDlkeme234

OpenStudy (anonymous):

are you joking? just bring across the 4.7 and add it to the -0.9 to get the answer

OpenStudy (anonymous):

right and the anwser is 5.6 right @BPDlkeme234

OpenStudy (anonymous):

you need to add 4.7 to -.9 which lessens the #

OpenStudy (anonymous):

no its no

OpenStudy (anonymous):

z =3.8

OpenStudy (anonymous):

bot

OpenStudy (anonymous):

not

OpenStudy (anonymous):

so 3.8 then

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

At noon the outside temperature was –5°F. By 4 p.m. the same day, the temperature had dropped 13°F. What was the outside temperature at 4 p.m.? At 4 p.m. the outside temperature was °F.

OpenStudy (anonymous):

jesus, I wish you took my tutorials in algebra pielover!123

OpenStudy (anonymous):

sorry lol

OpenStudy (anonymous):

slove for oF

OpenStudy (anonymous):

each side of the equation has to balance, so basically, we add 4.7 to each side of that equation

OpenStudy (anonymous):

did you look at the other problem?

OpenStudy (anonymous):

the -4.7 and the 4.7 you add on cancels, and the 4.7 you add to the Left hand side adds to the -0.9

OpenStudy (anonymous):

yea i know i got it right thank you :D

OpenStudy (anonymous):

but i need help with this other one @BPDlkeme234

OpenStudy (anonymous):

At noon the outside temperature was –5°F. By 4 p.m. the same day, the temperature had dropped 13°F. What was the outside temperature at 4 p.m.? At 4 p.m. the outside temperature was °F.

OpenStudy (anonymous):

slove for °F

OpenStudy (anonymous):

at 12am = -5F by 4pm = -5F (-13F) = -18F same question repeated

OpenStudy (anonymous):

oh sorry and thats what i got on paper

OpenStudy (anonymous):

Each Sunday, Mario weighs himself. Here are his weight changes over a 4-week period. Week 1 2 3 4 Weight Change –1.8 +0.9 –1.4 +0.4 How much has Mario's weight changed over the entire 4 weeks? A. 4.5 pounds B. 1.9 pounds C. –1.9 pounds D. –4.5 pounds

OpenStudy (anonymous):

@BPDlkeme234

OpenStudy (anonymous):

pls hurry

OpenStudy (anonymous):

@BPDlkeme234 do you know the answer???

OpenStudy (anonymous):

well he lost 1.8 in the first week, gained 0.9 in the second week, lost 1.4 in the third week and gained 0.4 in the fourth week. so: (-1.8)+ (-0.9)+(-1.4) +(+0.4) = (-1.8) +(-1.4)+ (+0.9)+ (+0.4)= ((-3.2) +(+1.3) so overall he lost 3.2 and gained 1.3 = 1.3 - 3.2 = -1.9 i.e total loss of 1.9

OpenStudy (anonymous):

–1.8 +0.9 –1.4 +0.4 =? just evaluate this and you'll get it

OpenStudy (anonymous):

i would get -1.9

OpenStudy (anonymous):

i.e the negative sign means he lost 1.9

OpenStudy (anonymous):

so it would be B then right

OpenStudy (anonymous):

correct answer = B

OpenStudy (anonymous):

thanks you so much :D

OpenStudy (anonymous):

sorry @pielover answer = C

OpenStudy (anonymous):

;/ i wish would have told me that i just turned in my test and got a 60%

OpenStudy (anonymous):

my apologies

OpenStudy (anonymous):

it's ok

OpenStudy (anonymous):

congrats by the way!

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

pielover, did you think I was arrogant?